A flashlight underwater (n=1.43) is aimed at the surface with incident angle 30°. What is the angle of refraction in the air (n=1) ?

1
Expert's answer
2022-03-07T11:44:02-0500

The Snell's law says

"n_i\sin\theta_i=n_r\sin\theta _r"

Thus

"\theta_r=\sin^{-1}\left(\frac{n_i}{n_r}\sin\theta_i\right)"

"\theta_r=\sin^{-1}\left(\frac{1.43}{1.00}\sin30^\circ\right)=45.6^\circ"